A Biologist mixes a 15% salt solution with a 35% salt solution to create 20 mL of a 22% salt solution.Which equation can you use to find the amount of 15% salt solution that is used?
Answer by ptaylor(2198)

Let x=amount of 15% salt solution that is used
Then 20-x=amount of 35% salt solution that is used
Now we know that the amount of pure salt that exists before the mixture takes place (0.15x+0.35(20-x)) has to equal the amount of pure salt that exists after the mixture takes place (0.22*20). Soooooo:
0.15x+0.35(20-x)=0.22*20 Now if we solve this equation for x, we get the amount of 15% salt solution that is used.
SO, LET'S DO IT!
Simplifying, we get:
0.15x+7-0.35x=4.4 subtract 7 from each side
-0.20x=-2.6
x=13 mL----amount of 15% salt solution that is used
CK
0.15*13+0.35*7=4.4
1.95+2.45=4.4
4.4=4.4
Hope this helps---ptaylor